Assume a member is selected at random from the population represented by the graph. Find the probability that the member seleA
Answer #1

Solution :

Given that, X ~ N(20.8, 5.2²)

μ = 20.8 and  σ = 5.2

The probability that the member selected at random is from the shaded area of the graph would be given by, P(26 < X < 32).

P(26 < X < 32) = P(X < 32) - P(X ≤ 26)

We know that if X ~ N(μ, σ²) then, X-M Z=1 ~ N(0,1) o

\large \therefore P(26 < X < 32) = P\left ( \frac{X-\mu}{\sigma} < \frac{32-\mu}{\sigma} \right ) - P\left ( \frac{X-\mu}{\sigma} \leq \frac{26-\mu}{\sigma} \right )

\large \therefore P(26 < X < 32) = P\left (Z < \frac{32-20.8}{5.2} \right ) - P\left (Z \leq \frac{26-20.8}{5.2} \right )

\large \therefore P(26 < X < 32) = P\left (Z < 2.1538 \right ) - P\left (Z \leq 1 \right )

Using "pnorm" function of R we get,

P(Z < 2.1538) = 0.9843 and P(Z ≤ 1) = 0.8413

\large \therefore P(26 < X < 32) = 0.9843 - 0.8413

\large \therefore P(26 < X < 32) = 0.1430

Hence, the probability that the member selected at random is from the shaded area of the graph is 0.1430.

Please rate the answer. Thank you.